PREX2RunPlan2019

From PREX Wiki
Jump to: navigation, search


CREX Main << PREX II Main

PREX Run Plan for   Commissioning   and   Auxiliary Measurements

See also the Daily Run Plan and Earlier Daily Run Plan and Auxiliary Measurements

PREX I - 2010 Run Plan


The experiment is approved for 35 PAC days including 25 production days. There are 49 days on our calendar.

Below is a list of tasks in the approximate time order, with person(s) responsible, approximate amount of time required, and comments about the conditions or needs. The time will, in most cases, be broken up into manageable chunks. For example, the GEM commissioning will occur in about 4 periods of time, each 4 hours, separated by at least one day. After commissioning we go into a regular production mode; we will provide an updated document webpage for shift worker instructions.

Total time: ?   Based on this list, each day we will come up with an adjustment to the Daily Run Plan. Some flexibility will be needed to accomodate problems and changes in plans.

Beam Restoration

  1. Basic beam setup 1 shift: Yves Roblin
    • establish tune beam and low current CW beam in the hall.
    1. Establish beam to dump, through Compton chicane
    2. Design match to hall expected as part of original configuration, if additional beam time is required will be necessary to coordinate with optics/production runplan
    3. ion chamber calibration will be skipped until we can get targets in (MASK ION CHAMBERS); CG will coordinate
  2. Source and Parity Quality Beam 0 shift (complete before 6/17): Caryn Palatchi, Amali Premathilake, Kent Paschke
    1. Source laser optics configuration and initial setpoints done
    2. Wein left/ Wein Right injector optics configuration done
    3. Characterize injector optics configuration done


Low Current Commissioning

  • Basic beamline commissioning (may be 3 shifts total)
    • Establish low-current beam with spectrometer magnets on. Establish centering on collimator. BPM offset calibration. Commission target ladders.
    • Preparations:
      • Stripchart on ion chambers, with bpm4e and bpm4a positions
      • stripchart on collimator and beamline temps with bpm 4a and 4e positions
      • stripchart on compton rates (with beam current) no more 2x what is present in HACOMPTONLOG
  1. Verify Beamline setup 1 shift Yves Roblin, Bob
    • Tune mode. No target. Raster off. Ion chambers masked. ATLis 19326
    1. beam transport through Compton chicane up to 5uA CW.
      • Rates in finger scintillators no more 2x what is present in HACOMPTONLOG
      • ComptonUS1 ~ 30 Hz, ComptonDSbg1 ~ 870 Hz, ComptonDCbg2 ~ 1010 Hz
    2. spot size at target (harps) (nominal spec is 150 um x 150 um. Larger is better)
      • Use 5uA CW, but keep eye on compton rate, ion chambers, temps.)
    3. Tune beam. Turn on septum and verify still on dump. Clean transport to dump for PREX set points
    4. 1 arm running (Q1 on) test to verify no beam motion.
    5. Turn on both Q1 spectrometers.
  2. Hall bullseye scan 0.5 shift Bob
    • harp scans at each of 5 points around nominal center, (0,0), (2,2), (2,-2), (-2,-2), (-2,2).
    1. Set prescale for high rate of clock triggers
    2. Watch beam on viewer, ion chamber and temperature stripcharts.
    3. use 2-5uA CW to find each corner position. 4A, 4E harp scan for one of the corner positions and verify EPICS readback of BPMs, check spectrometer BPM calibrations.
      1. If PREX expert is available, parasitically collect data for low current beam monitoring (cavities). If during day have RC inform J Musson before taking this data.
  3. Collimator checkout 0.5 shift Yves, Ciprian G, Kent P
    • will verify that we are cleanly going through the collimator and cross check power deposition calculations
    1. Target ion chamber setpoints can be set high to allow this test.
    2. setup ion chamber strip chart
    3. Establish 2-5uA tune mode, with empty target. Use 4x4 raster.
    4. move beam in 2mm steps as much as +/- 1 cm in both horizontal and vertical. Watch for rise the ion chamber signal. Record positions where edges of the collimator bore are found.
    5. set beam to center on colllimator based on those edges.
    6. compare to position of carbon hole on warm and cold ladders. Do we need to recheck target frames for targets? Do we need to change encoder positions?
  4. Warm target position 0.25 shift Kent (Silviu, present or on call?)
    • Verify target alignment.
    1. insert warm ladder, carbon hole.
    2. Set raster for 8mmx8mm
    3. establish 2uA CW. Run spot++ to find carbon hole. log it.
    4. Adjust raster setting to establish 4x5 raster size. Log it.
    5. if carbon hole not in 4x4 raster view, need to reconsider. (move beam, or use off center?)
  5. Cold target position 0.25 shift Kent (Silviu, present or on call?)
    • Verify target alignment.
    • target and dump ion chamber set points will need to be raised to perform this commissioning
    1. insert warm ladder, carbon hole.
    2. Set raster nominal 8x8 (or what could be used.)
    3. establish 5uA CW. Run spot++ to find carbon hole. Log it.
    4. Set raster to 4x5 set point from before. Use spot++ to verify it.
    5. if carbon hole not in 4x5 raster view, need to reconsider. (move beam, or use off center?)
    6. Move to lead-208 target, location #1(?)
    7. establish 5uA CW, 4x5 raster. Run spot++ to verify raster size, not hitting edges. Log it.
  6. Ion Chamber calibration - thick Target 0.25 shift Ciprian G
    • Procedure discussed with Eric Forman: thick target calibration procedure (ramp 2-10uA) for ALL ion chambers (including the dump)
    • Must keep rates in compton finger scintillators; no more 2x what is present in [https://logbooks.jlab.org/entry/3687977
    1. Establish beam, 5uA beam, 4x4 raster, on lead target.
    2. move beam to x=+1mm (relative to collimator center)
    3. target and dump ion chamber setpoints can be set high this configuration
    4. Execute thick target Ion Chamber calibration, From Atlis 19326. This calibration will be used to set ion chamber setpoints.
      1. Dump and Near-target ICs: set for 45 degree targets by projecting ion chamber setpoints for 20uA current or 90 degree targets by projecting to 70uA current.
    • parasitic: first check of SAMs at low current (call Devi? or Dustin? before starting)
  7. Very low current commissioning 0.25 shift Yves. Ye Tian. Caryn.
    • commission low current cavities relative to stripline calibration. Commission cavity bpm lock.
    • Note: position scan and lock setup are separate pieces, can happen at different times.
    • Either/ both can happen later (during optics running). Needed before sieve-out Q2 studies.
    1. Cavity lock
      • [Yves is required for step]
      1. setup cavity stripchart. Establish 2uA CW.
      2. setup cavity slow lock. Verify that in operation this lock holds the position in the cavity epics readback.
      3. lower current to 100 nA. Verify that the lock still holds the cavity epics readback.
    2. Cavity Position scan.
      1. setup cavity stripchart. Establish 1uA CW.
      2. Find corrector positions for +/-2mm in X, Y.
      3. Reduce current to 50nA.
      4. set to each corrector position, verify position measurements with epics readback
        1. Turn on both parity DAQ and counting DAQs. For the counting DAQ, the trigger is set to T1=1, and make sure that S0 HV is on.
          1. For changing the trigger prescale factor: for the LHRS, please go to adaq@adaq3, type prescaleL, and you will see a gui, and set T1=1 and press the save button and exit it.
          2. For the RHRS, please go to the adaq@adaq2, typle prescaleR, and you will see the same gui, just set T1=1 too and press the save button and exit it.
        2. Start/and stop the parity DAQ and counting DAQs at same time for the each position point.
        3. Look at the strip chart of IPM1H04BX, IPM1H04CX, IPM1H04DX and IPM1H04BY, IPM1H04CY, (current signal) IPM1H04DY, IPM1H04B, IPM1H04C, IPM1H04D, if the signal is saturate or too small to present. Then we need to change the gain of the cavity BPMs
          1. Change the gain, please call Ye Tian (803-553-3570)

Beam Monitor calibration

  1. Establish 50 uA in hall
  2. Run current scan, local around 50uA, using Unser to calibrate beam charge monitor readout in parity DAQ.
    1. Run the Compton DAQ concurrently, for calibration
    2. Go to FixedGain when at 50 uA on bpms (list of crates???) and repeat local current scan to set bpm pedestals.

Spectrometer Commissioning

Detailed Optics Plan Name and time assignments needs to be worked out by optics crew

  1. Tracking checkout 1 shift: Bob M, Ryan R, Chandan G
    1. Sieve In, raster off. Tune B. 45 deg C target. 1 uA.
    2. Septum 377 A
    3. Check rates with scalars.
    4. Set prescales for S0 triggers.
    5. VDC checkout; look at tracks, VDC spectra
    6. Some short run for GEM: noise, pulse-heights, tracks, etc. After this GEMS out of DAQ to improve efficiency.
  2. Sieve reconstruction verification 0.25 shift:Seamus or Bob, Ryan
    1. Sieve In, raster off. Tune B. 45 deg C target. 1 uA.
    2. Septum 377 A
    3. 500k events
  3. Tune septum to central ray 1 shift: Seamus, Ryan
    1. Sieve In, raster off. Tune B. 45 deg C target. 1 uA.
    2. Beam centered on axis
    3. Start septum at 377 A
    4. 500k events per run
    5. Adjust Q1 fields by 10% low (one run) and high (one run)
      1. Identify the invariant position in the phitarg-thetaTarg space and adjust septum to move the central ray to expected location
    6. Repeat 10% low (one run) and high (one run) to verify central hole in correct location
    7. Verify for Q2 and Q3 that central hole does not move under field changes (1 run each, 10% low). (optional)
      • Expected: the hole will be invariant under all quad changes. This location is about 2mm toward beam from central hole.
    8. Last step: CYCLE QUADS to get back to precise Tune B optics
  4. Inner Edge Identification 0.5 shift: Seamus, Ryan
    1. Sieve In, raster off. Tune B. 45 deg C target. 1 uA.
    2. Beam centered on axis
    3. Start septum at central tune
    4. 500 k events per run
    5. Take 9 points over septum current from +10% to -10%
      1. The expected inner edge of the acceptance should match simulation
      2. If there is a small-angle obstruction for currents at or above the central tune current, identify the lower septum current where there is no inner obstruction for both arms. Use this as new production septum current.
    6. Last step: Cycle Septum
  5. Q1 acceptance scan 0.25 shift:Ye T
    1. Sieve In, raster off. Tune B. 45 deg C target. 1 uA.
    2. Beam centered on axis
    3. Optics at central tune
    4. 500 k events per run
    5. Take a run for Q1 filed +/-10% over 5 points (cycle, then scan top to bottom)
      • Expected: later analysis should show out of place acceptance compares to simulation.
  6. Quartz spot width minimization 0.5 shift: Ryan, Seamus? Bob?
    1. Sieve IN (If this is last step of optics configuration, take an access to pull sieve OUT)
    2. raster ON (4x6 raster)
    3. Start with central tune
    4. 45 deg C target. 1 uA.
    5. Beam centered on axis
    6. Optics at central tune
    7. 500 k events per run
    8. Take a run along eigenvector quads +/-10% for Q1 filed +/-10% over 5 points (cycle, then scan top to bottom)
      • Expected: width should be minimized
      • Tune that minimizes width at detectors is new central tune.
  7. Quartz detector checkout 1 shift: Dustin M, Devi A., VDC expert
    1. Sieve OUT
    2. raster ON (4x6 raster)
    3. Start with central tune
    4. thin C target. 0.5uA - 1 uA (want quartz scaler rates below 1MHz but need beam position monitoring)
    5. Beam centered on axis
    6. Optics at (new) central tune
    7. verify small spot at quartz (thin C, low current, sieve out, raster on?)
    8. thin C, low current, raster on, Sieve OUT: positions Plan
    9. PMT spectra: Examine pulse height distributions to get RMS/mean and PEyields
    10. rates: These will come from scalers. We can't do a rate "check" without going to integrate mode. This will be done during Width study
  8. Optics calibration data 0.5 shift Nilanga, Siyu
    1. Sieve In, raster off. Tune B. 1 uA.
    2. 45deg Tungsten target
    3. Beam centered on axis
    4. Optics at central tune
    5. 500 k events per run
    6. spectrometer momentum scan (details in run plan still)
  9. Water cell pointing 1 shift Nilanga, Siyu
    1. Sieve IN: pointing measurement
    2. Sieve OUT: pointing measurement
    3. (details in run plan still)
  10. Y target reconstruction on warm and cold ladders
    1. Raster off.
    2. 45 deg carbon
    3. 90 deg carbon
  11. Q2 measurement 0.5 shift:
    1. Production target
  12. Quartz detector rate estimate 0.5 shift KK, Dustin, Devi, Tao
    1. controlled access to change UpStream quartz detector (in each arm) to integrating mode
    2. draft plan
  13. Position At detectors : Dustin, Ryan

High current checkout

detailed plan

  1. Basic setup 0.5 shift Yves R, Ciprian G:
    • No target.
    • 50uA beam current, through Compton (backgrounds on fingers and Compton photon detector).
  2. High current with target 0.5 shift Ciprian G, Dustin M:
    1. Establish 4x4 raster. Insert lead target.
    2. Detector and SAM voltages (Dustin M)
  3. Beamline instrumentation calibration 0.5 shift Caryn P,
    1. Current ramp for BPM/BCM calibrations and detector pedestals detailed plan
  4. Initial production condition test 0.25 shift Caryn P, detailed plan
    1. detector and SAM widths
    2. regression sensitivities
    3. beam noise and asymmetry check
    4. collimator water and thermocouple temperatures,
    5. radiation monitors and ion chamber stability
  5. Width studies 0.5 shift KK,
    1. detector and SAM widths, targets, shutters, etc
    2. deinstall shutters at conclusion of test

Establish Production Conditions

  1. Beam modulation 0.5 shiftVictoria O., Ye T,: (Detectors needed for BeamMod commissioning)
    1. turn on
      1. In integrating mode
      2. check HV of Main Detector
      3. Lumis on
    2. set Amplitudes
      1. start with .100 A set amplitude. Run (at least) 2 supercycles. Check online gui plots for response.
      2. increase amplitudes by .50 A, to see response (if needed).
    3. test orthogonality
      1. after minimum 2 good supercycles (more is better) ready to examine sensitivities
      2. if orthogonality is bad, need additional time to discuss with experts about retuning beam.
    4. turn off
      1. restore conditions for Moller commissioning
  2. Detector checkout 1 shift Dustin M:
    1. quartz detector position: verify no inelastic event acceptance (plan needed)
    2. width study (plan needed)
    3. raster sync (plan needed)
  3. Spin Dance 1 shift: Simona Malace, Sanghwa Park
    1. Moller polarimeter commissioning (establish beam) (plan needed)
    2. Short spin dance (run plan needed, established optimal spin angle for this Wien state) (plan needed)
  4. Hall PQB 1 shift Caryn P,:
    1. Check asymmetries in the hall
    2. Adiabatic dampening / Matching (plan needed)
      1. Use helicity magnets to tune phase trombone
    3. Establish feedback (if Hall C is not on schedule their feedback for later)
  5. Production: 0.5-3 shift? All
  6. AT measurement: 1.5 shift? All (plan needed)
    1. Do we need to verify zero longitudinal using Moller polarimeter?
    2. Carbon
    3. Lead
    4. Calcium
  7. Initial Moller commissioning 1 shift Simona M, Sanghwa P,
    1. minimal set for first absolute measurements
  8. Initial Compton commissioning 1 shift Dave G, Juan Carlos C,
    1. minimal set for measurement conditions

Start Production

Important Activities After First Few Days

  1. Linearity Studies --
  2. Background Studies --
    1. Scans of Septum Magnet and HRS Dipole
    2. Thin Lead Target to check for inelastics at high-resolution
    3. Thin C12 to measure diamond background
  3. Possible scan of Q1 to optimize acceptance and verify that collimator defines acceptance. (old idea, still relevant?)
  4. We will start with the Wein in Spin-Right. Once we switch to Spin-Left we need to do another spin-dance.

Polarimetry Studies

  1. Compton Commissioning --   Gaskell, et.al.
    1. Beam Tune, Background reduction -- Bteam,
    2. Compton Cavity Checkout --
    3. Photon Detector Checkout -- CMU group
  2. Moller Commissioning -- Simona Malace, Don Jones, et.al.
    1. Magnet Alignment -- 1 shift (swing)
    2. Raster size and pulse-mode -- 1 shift (day)
    3. Target commissioning - 3 shifts
    4. Pulse-mode target commissioning - 2 shifts
    5. DAQ checkout -- 1 shift


Important Activities During run

  1. Linearity Studies --
  2. Background Studies --
  3. Repeat Q2 and pointing measurements
  4. Arc Energy Measurement